# Domain Generalization for Medical Image Analysis: A Survey

### v2

@@ 1 Introduction
[…]
With homogeneous data distribution, well-designed models perform on par with and often surpass their human counterparts in many applications. However, their reliability and robustness can be compromised when presented with previously unseen, out-of-distribution, or heterogeneous data. This highlights a common challenge in the field of MedIA: the limited capacity of models to generalize to unfamiliar data distributions. Changes in data distribution can result from variations in imaging equipment, protocols, or patient populations. Domain generalization aims to overcome these challenges by developing models that can adapt to new, unseen domains without compromising performance.
1.1 Domain Generalization for Medical Image Analysis
@@
  • Image appearance variability: Variability in medical imaging refer to differences and inconsistencies typically manifest during the data acquisition process [4]. These variability may arise externally from using different modalities, protocols, scanner types, and patient populations across multiple healthcare facilities, while internal variability may also occur within a controlled setting (e.g., same scanner or healthcare facility) due to factors such as hardware aging, software parameter variations, and human error (e.g., human motion).
@@
1.2 Our Contributions
With these challenging factors in mind, this survey provides a comprehensive review of domain generalization techniques specifically tailored to MedIA. There already exist few survey papers on domain generalization with a specific focus on MedIA, but these are focused on limited data domain and task, i.e., mammography-based mass detection [7], electroencephalography (EEG)-based emotion assessment [8], and computational pathology [9]. Also, there are several survey papers on related topics for MedIA, such as domain adaptation [10, 11] and harmonization [4, 12]. However, domain generalization presents unique challenges compared to these tasks.
@@ 1.2 Our Contributions
Our survey aims to provide a holistic view of how domain generalization techniques interact within the broader structure of a MedIA system. We go beyond the methodological hierarchy presented in previous surveys and delve into the operational consequences of domain generalization on the entire MedIA workflow (see Fig. 1). Our focus is on understanding how domain generalization can be seamlessly integrated into every step of the decision-making process, including but not limited to data acquisition, pre-processing, model prediction, and analysis. To this end, we categorize domain generalization techniques into each step of the MedIA workflow, i.e., from data preparation to analysis.
2 Background
2.1 Problem Definition
@@
2.2 Settings of Domain Generalization
This subsection elucidates different settings for DG in the MedIA workflow, detailing the various configurations and challenges present in both the source and target domains during the implementation process (see Fig. 2).
2.2.1 Settings for Source Domain
DG typically focuses on two settings regarding the number of source domains: multi-source DG and single-source DG [14]. The multi-source setting assumes multiple distinct but relevant domains are available (i.e., $M>1$). By leveraging the data from these domains, representations invariant to disparate marginal distributions are learned. This is usually accomplished by minimizing the domain discrepancy among the source domains during the training process. The single-source setting assumes training data is homogeneous (i.e., $M=1$). Therefore, this setting does not require domain labels during training. Single-source DG tends to be more challenging than multi-source DG as it may not capture the full diversity of data distributions that exist across different domains. Refer to Section 4 for some extreme settings for domain generalization (such as, open-set DG, source-free DG, and unsupervised DG).
2.2.2 Settings for Target Domain
@@
  • Cross-site DG: As the most prevalent form of DG for MedIA, the goal of cross-site DG is to develop models that can generalize well across different medical imaging datasets collected from multiple healthcare institutions. Cross-site DG helps in creating more robust models that can be deployed across different healthcare institutions without the need for extensive site-specific fine-tuning.
@@
2.2.3 Settings for Domain Shift
@@
2.3 Related Machine Learning Tasks
TABLE II: Related machine learning tasks categorized by covariate shift, concept shift, and access to a target domain $\mathcal{S}_{target}$.
@@ 2.3 Related Machine Learning Tasks
  • Out-of-distribution (OOD) Generalization aims to detect the concept shift between in-distribution (ID) and OOD data. While OOD and DG both assume no access to the target domain, the main difference between OOD and DG lies in that they focus on different domain shifts. Specifically, OOD mainly focuses on concept shift whereas DG considers both covariate and concept shift in their problem settings.
  • Zero-shot Learning (ZSL) is closely related to OOD generalization in that it aims to classify test samples with concept shift, but ZSL generally leverages auxiliary information, such as attribute descriptions, related to the target domain.
@@
{forest}
[…]
3 Methods
In this section, we review and explain a series of DG methods for medical imaging. We employ a bottom-up approach and categorize the methods into data-level, feature-level, model-level, and analysis-level DG methods (see Fig. 3). Then, we explore some DG methods under extreme constraints. Readers are referred to Appendix C in the Supplementary for a summary of literature, critical analysis, and pratical suggestion for MedIA pipeline.
@@ 3 Methods
3.1 Data-level Generalization
@@ 3.1.1 Data Manipulation
3.1.1.1   Image Processing Methods
Image processing techniques involve closed-form or learnable transformation functions to increase the diversity and quantity of training data. Examples of some traditional image processing methods include registration, resampling, and filtering, which are specifically designed for the distinct characteristics of the medical image data in question. Although many traditional image processing methods have empirically shown to improve the model’s generalizability [24], they are predominantly employed as pre-processing tools for downstream tasks, rather than as standalone solutions for DG. Also, with the advancement of deep learning, there has been a gradual shift towards incorporating these techniques directly into deep learning architectures, enabling a more seamless integration of end-to-end learning of domain-invariant features (see Section 3.2). Readers are referred to [25] for a comprehensive review of the traditional image processing methods. In the following paragraphs, we explore several deep learning-based image processing methods specifically designed for DG for MedIA tasks.
@@ 3.1.1.1   Image Processing Methods
3.1.1.2   Surrogate Methods
@@
3.1.2 Data Augmentation
[…]
3.1.2.1   Randomization-based Augmentation
The idea of random augmentation is to generate novel input data by applying random transformations to the original data. Some conventional techniques include randomly applying flipping, rotation, scaling, cropping, adding noise, etc., which are used extensively to improve a model’s generalization performance by reducing overfitting [54]. Li et al[55] developed a novel style transfer network that augments a domain by modifying cardiac images with randomly sampled shape and spatial (i.e., slice index) priors to alleviate the modality-level difference for cardiac segmentation. Liu et al[44] proposed a random amplitude mixup method that randomly mixes the amplitudes of different images for DG for fundus image restoration.
3.1.2.2   Adversarial-based Augmentation
Adversarial-based data augmentation methods operate on the principle of creating adversarial examples that aim to maximize the model’s uncertainty, thereby improving its robustness and generalizability. In this section, we concentrate on data-level adversarial augmentation, while a discussion on model-level adversarial training can be found in Section 3.3.1.3. Tomar et al[56] developed a method that combines knowledge distillation with adversarial-based data augmentation for cross-site medical image segmentation tasks. The process involves the creation of augmented data that is adversarial to the current model, to push the model’s feature representations toward the decision boundary. This is achieved by optimizing and sampling data augmentations that simulate data in the uncertain region of the feature space, thereby improving the model’s ability to generalize from the training data to unseen test data.
3.1.2.3   Generative Models
@@
3.1.3 Problem-specific Data-level Methods
@@
3.1.3.1   Cross-modal Generative Models
Cross-modal generative models represent a pioneering paradigm for achieving DG, wherein models are trained to gain knowledge of the data distribution across diverse modalities (e.g., CT, MRI, X-ray, and PET). These models, often based on Generative Adversarial Networks (GANs), generate synthetic data [55] or suitable latent representations [60], which bridge the distributional gap among cross-modalities. This strategy allows us to provide a model especially capable in medical imaging where data could vary greatly due to patient cohorts, hospital practices, or different imaging modalities. As obvious advantages of such a model, it can be highly valuable when one modality is unavailable for a particular patient or when the model is required to generalize to an unseen domain where a different imaging modality is used. Readers are refer to Xie et al[61] for a comprehensive review on cross-modal neuroimage synthesis.
@@ 3.1.3.2   Stain normalization
3.2 Feature-level Generalization
TABLE III: Summary of feature-level domain generalization methods.

### v3

# Domain Generalization for Medical Image Analysis: A Review

@@ I Introduction
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Data Acq. Image Recon. Upstream Down- stream Analysis Data-level §$\S$§III-AFeature-level §$\S$§III-BModel-level §$\S$§III-CAnalysis-level §$\S$§III-DAcq. standards [4]Image processing (p. III-A1a) Alignment (p. III-B1) Learning strategy (p. III-C1)Transferability (p. III-D1a)Frequency-based (p. III-A1b)Data augmentation (p. III-A2) Optimization (p. III-C3b) Causality (p. III-D2)Interpretable AI (p. III-D1)MedIA WorkflowDG Techniques
Figure 1: Overview of the medical imaging analysis (MedIA) pipeline illustrating various stages and their associated domain generalization (DG) techniques. Stages include data acquisition, image reconstruction, upstream feature extraction, downstream task, and analysis. Each stage is associated with references to specific sections (§$\S$), pages (p.), or external citations where the techniques are detailed.
With homogeneous data distributions, well-designed models perform on par with and often surpass their human counterparts in many applications. However, their reliability and robustness can be compromised when presented with previously unseen, out-of-distribution, or heterogeneous data. This highlights a common challenge in the field of MedIA: the limited capacity of models to generalize to unfamiliar data distributions. Changes in data distribution can result from variations in imaging equipment, protocols, or patient populations. Domain generalization aims to overcome these challenges by developing models that can adapt to new, unseen domains without compromising performance.
I-A Domain Generalization for Medical Image Analysis
@@
  • Image appearance variability: Variability in medical imaging refers to differences and inconsistencies typically manifest during the data acquisition process [4]. This variability may arise externally from using different modalities, protocols, scanner types, and patient populations across multiple healthcare facilities, while internal variability may also occur within a controlled setting (e.g., same scanner or healthcare facility) due to factors such as hardware aging, software parameter variations, and human error (e.g., human motion).
@@
I-B Our Contributions
With these challenging factors in mind, this review provides a comprehensive review of domain generalization techniques specifically tailored to MedIA. There already exist few review papers on domain generalization with a specific focus on MedIA, but these are limited to specific data domain and task, i.e., mammography-based mass detection [7], electroencephalography (EEG)-based emotion assessment [8], and computational pathology [9]. Also, there are several review papers on related topics for MedIA, such as domain adaptation [10, 11] and harmonization [4]. However, domain generalization presents unique challenges compared to these tasks.
@@ I-B Our Contributions
Our review aims to provide a holistic view of how domain generalization techniques interact within the broader structure of a MedIA system. We go beyond the methodological hierarchy presented in previous surveys and delve into the operational consequences of domain generalization on the entire MedIA workflow (see Fig. 1). Our focus is on understanding how domain generalization can be seamlessly integrated into every step of the decision-making process, including but not limited to data acquisition, pre-processing, model prediction, and analysis. To this end, we categorize domain generalization techniques into each step of the MedIA workflow, i.e., from data preparation to analysis.
I-C Scope of Review
Figure 2: Number of publications per year on the Google Scholar database.
Literature search and selection were conducted by researchers experienced in machine learning and medical image analysis. We used the Google Scholar search engine with three different search strategies, resulting in a database of 1,621 papers. First, we searched for all papers that cited the existing domain generalization review [12, 13, 14, 19, 20, 11, 8]. Second, we searched Google Scholar using the exact phrase “domain generaliz(s)ation” and medical-related keywords (“Medical, CT, ultrasound, MRI, PET, X-ray, histology, histopathology, pathology, fundus, dermoscopy, endoscopy, mammography”) and selected 1,000 papers sorted by relevancy. Lastly, we searched for the top 1,000 papers with the terms “unseen”, “domain”, and medical-related keywords. The eligibility criteria for papers to be included in this review are that they have conducted at least one experiment involving the use of medical images within the domain generalization problem settings (see §$\S$§II-A), regardless of their use of the term “domain generalization” in their paper (small number of papers instead use “unseen”, “out-of-distribution”, or their own terms). Peer-reviewed published papers were prioritized, but non-peer-reviewed archive papers (e.g., arXiv, bioRxiv) were also included if they had been deemed particularly suitable for selection (e.g., highly relevant, highly significant, highly cited).
II Background
@@ II-A Problem Definition
II-B Settings of Domain Generalization
This subsection elucidates different settings for DG in the MedIA workflow, detailing the various configurations and challenges present in both the source and target domains during the implementation process (see Fig. 3). A critical aspect of DG is overcoming the domain shift — differences between the source and target domains that hinder the model’s ability to generalize. These gaps typically arise due to variations such as:
  • Intensity Variations: Differences in image brightness or contrast, e.g., varying exposure levels in chest X-rays or staining intensities in histology images.
  • Resolution Differences: Variations in image detail, such as differences between high and low-resolution ultrasound devices or varying pixel densities in fundus photographs.
  • Noise Characteristics: Varying levels of image noise, e.g., speckle noise in ultrasound or grain in low-dose CT scans.
  • Artifact Patterns: Modality-specific artifacts, such as motion blur in MRI, beam hardening in CT, or light reflections in endoscopy images.
  • Anatomical Variations: Differences in patient populations, leading to variations in organ sizes or shapes across different datasets.
  • Label Distributions: Varying disease prevalence or severity between domains, affecting class balance in tasks like skin lesion classification from dermatology images.
  • Acquisition Protocols: Differences in imaging techniques, such as varying MRI sequences, CT reconstruction kernels, or staining protocols in histopathology.
As it is often challenging to attribute domain shifts to a single cause, we have systematically categorized these variations with into cross-site, cross-modality, cross-temporal, and cross-protocol domain shifts (see §$\S$§II-B2). These shifts can be further classified into covariate and concept shifts (see §$\S$§II-B3), providing a comprehensive framework for understanding and addressing domain shifts in MedIA.
II-B1 Settings for Source Domain
DG typically focuses on two settings regarding the number of source domains: multi-source DG and single-source DG [13]. The multi-source setting assumes multiple distinct but relevant domains are available (i.e., $M>1$). By leveraging the data from these domains, representations invariant to disparate marginal distributions are learned. This is usually accomplished by minimizing the domain discrepancy among the source domains during the training process. The single-source setting assumes training data is homogeneous (i.e., $M=1$). Therefore, this setting does not require domain labels during training. Single-source DG tends to be more challenging than multi-source DG as it may not capture the full diversity of data distributions that exist across different domains. Refer to Section IV for some extreme settings for domain generalization (such as, open-set DG, source-free DG, and unsupervised DG).
II-B2 Settings for Target Domain
@@
  • Cross-site DG: As the most prevalent form of DG for MedIA, the goal of cross-site DG is to develop models that can generalize well across different medical imaging datasets collected from multiple healthcare institutions. Cross-site DG helps in creating more robust models that can be deployed across different healthcare institutions without the need for extensive site-specific fine-tuning. Cross-site DG research often define a domain or ‘site’ as an individual healthcare facility, but some define it as an individual vendor, scanner, or even an individual patient when inter-patient variability is extremely large, which is a common setting in many EEG studies.
@@
II-B3 Settings for Domain Shift
@@
{forest}
for tree = draw=none, rounded corners, font=, text width=9em, text centered, calign=edge midpoint, edge = -Stealth, if level = 0fill=teal!30, text width=15em, if level¿= 1grow’=0, folder, folder indent=4mm, l sep=7mm, s sep=2mm, fill=teal!10, minimum height=2.5em, , if level = 1fill=teal!20, minimum height=0,, [Domain Generalization for MedIA [Data-level §$\S$§III-A [Manipulation §$\S$§III-A1] [Augmentation §$\S$§III-A2] [Problem-specific §$\S$§III-A3] ] [Feature-level §$\S$§III-B [Alignment §$\S$§III-B1] [Disentanglement §$\S$§III-B2] [Others §$\S$§III-B3] ] [Model-level §$\S$§III-C [Learning Strategies §$\S$§III-C1] [Model Framework §$\S$§III-C2] [Others §$\S$§III-C3] ] [Analysis-level §$\S$§III-D [Interpretable AI §$\S$§III-D1 [Transferability (p. III-D1a)]] [Causality §$\S$§III-D2] ] ]
II-C Related Machine Learning Tasks
TABLE II: Related machine learning tasks categorized by covariate and concept shift, and access to $\mathcal{S}_{target}$.
@@ II-C Related Machine Learning Tasks
  • Zero-shot Learning (ZSL) is closely related to out-of-distribution (OOD) generalization in that it aims to classify test samples with concept shift, but ZSL generally leverages auxiliary information, such as attribute descriptions, related to the target domain.
@@
  • OOD Generalization aims to detect the concept shift between in-distribution (ID) and OOD data. While OOD and DG both assume no access to the target domain, the main difference between OOD and DG lies in that they focus on different domain shifts. Specifically, OOD mainly focuses on concept shift whereas DG considers both covariate and concept shift in their problem settings.
III Methods
In this section, we review and explain a series of DG methods for medical imaging. We employ a bottom-up approach and categorize the methods into data-level, feature-level, model-level, and analysis-level DG methods (see Fig. 4). Then, we explore some DG methods under extreme constraints (see §$\S$§IV).
@@ III Methods
This categorization is inspired by the various stages of the MedIA pipeline, i.e., stages starting from data acquisition, image reconstruction, feature extraction, downstream task to analysis (see Fig. 1). Our aim is to provide researchers and practitioners a holistic view of how different approaches can be combined and applied throughout the MedIA pipeline.
TABLE III: DG techniques applicable to multiple levels of the MedIA pipeline.
Data-level Feature-level Model-level
Augmentation §$\S$§III-A2 §$\S$§III-B3a
Adversarial §$\S$§III-A2b §$\S$§III-C1c
Contrastive §$\S$§III-B2a §$\S$§III-C1b
It is important to note that some DG techniques can be applied at multiple levels of the MedIA pipeline (see Table III). Augmentation at the data-level (§$\S$§III-A2) involves directly manipulating input images to increase data diversity, while feature-level augmentation (§$\S$§III-B3a) operates on extracted feature representations to expand the feature space. Adversarial techniques at the data-level (§$\S$§III-A2b) focus on generating challenging examples to improve robustness, whereas at the model-level (§$\S$§III-C1c) they aim to learn domain-invariant features through adversarial training. Contrastive learning at the feature-level (§$\S$§III-B2a) encourages similar samples to have close representations while pushing dissimilar ones apart, while at the model-level (§$\S$§III-C1b) it is used as a self-supervised pretraining strategy to learn generalizable representations. These multi-level applications of similar concepts highlight how DG techniques can be integrated throughout the pipeline, each tailored to address domain shift at different stages of processing.
III-A Data-level Generalization
TABLE IV: Data-level domain generalization methods. Methods categorized by different settings for source and target domains (see §$\S$§II-B), task, organ, and modality used in experiment.
Method Specific Ref. Source Target Task Organ Modality
Image Processing Intensity Normalization [22] Multiple Site Segmentation Brain MRI
[23] Single Site, Sequence Segmentation Brain, Prostate, Cardiac MRI
Histogram Matching [24] Multiple Site Segmentation Cardiac MRI
[25] Multiple Site Segmentation Atrial MRI
[26] Multiple Site Segmentation Retinal Fundus
Color Normalization [27] Multiple Site Detection Tissue Histology
[28] Multiple Site Classification Retinal Fundus
[29] Multiple Site Classification Skin Dermatology
[30] Single, Multi Site Segmentation, Classification Retinal, Chest Fundus, X-ray
Surrogate Frequency-based DG [31] Multiple Site Segmentation Brain MRI
[32] Single, Multiple Site Segmentation Retinal Fundus
[33] Multiple Site Restoration Retinal Fundus
[34] Multiple Site Segmentation Retinal Fundus, OCT
[35] Single Site Segmentation Retinal Fundus
Using Raw Signals [36] Multiple Site Segmentation Brain MRI
[37] Single Sequence Segmentation Brain MRI
Dictionary Learning [38] Multiple Sequence Reconstruction Brain MRI
[39] Single Site Segmentation Prostate, Retinal MRI, Fundus
Augmentation Randomization-based [40] Single Sequence Segmentation Cardiac MRI
[33] Multiple Site Restoration Retinal Fundus
[41] Single Site Segmentation Retinal Fundus
[42] Single Site Segmentation Retinal Fundus
Adversarial-based [43] Multiple Site Segmentation Spinal, Prostate, Colon MRI, Histology
Generative Model [44] Multiple Site Classification Colon, Breast Histology
[45] Single Site Classification Colon Histology
[46] Single Sequence, Site Detection Liver PET, CT
Problem-specific Cross-modal Generative Model [47] Multiple Modality, Sequence Segmentation Brain, Prostate, Abdominal MRI, CT
[48] Single Modality, Sequence Segmantation Prostate, Abdominal, Cardiac MRI, CT
[49] Single Modality Segmentation Cardiac, Abdominal MRI, CT
Stain Normalization [50] Multiple Site Detection Breast Histology
[51] Multiple Site Classification, Segmentation Tissue, Breast Histology
@@ III-A1 Data Manipulation
III-A1a   Image Processing Methods
Image processing techniques involve closed-form or learnable transformation functions to increase the diversity and quantity of training data. Examples of some traditional image processing methods include registration, resampling, and filtering, which are specifically designed for the distinct characteristics of the medical image data in question. Although many traditional image processing methods have empirically shown to improve the model’s generalizability [53], they are predominantly employed as pre-processing tools for downstream tasks, rather than as standalone solutions for DG. Also, with the advancement of deep learning, there has been a gradual shift towards incorporating these techniques directly into deep learning architectures, enabling a more seamless integration of end-to-end learning of domain-invariant features (see Section III-B). Readers are referred to [54] for a comprehensive review of the traditional image processing methods. In the following paragraphs, we explore several deep learning-based image processing methods specifically designed for DG for MedIA tasks.
@@
It is worth noting that some harmonization techniques, while primarily aimed at reducing non-biological variability across imaging sites or protocols, can be applied in a DG setting. A notable example is the ComBat harmonization [61], which uses an empirical Bayes framework to adjust for site effects while preserving biological variability. More recently, deep learning-based harmonization techniques have emerged. Those include autoencoders with heavy regularization or normalization layers [62, 63], and generative adversarial networks (GANs) capable of scanner-to-scanner translation [64, 65]. Some of those methods, such as adversarial learning with domain classifiers [66, 67] and conditional variational autoencoders [68], showed the capability of deriving scanner-/cohort-invariant features for reconstructing harmonized samples across unseen scanners.
III-A1b   Surrogate Methods
@@
III-A2 Data Augmentation
Data augmentation is one of the most prevalent and important forms of DG in MedIA. It refers to techniques that artificially expand and diversify the training dataset by applying various transformations to existing data. The primary goal is to improve the model’s ability to generalize across different domains by exposing it to a wider range of data variations during training. Unlike feature-level augmentation (§$\S$§III-B3a), which modifies the learned feature representations, data-level augmentation directly alters the input data-space.
The widespread adoption and significance of data augmentation in DG for MedIA stem from its effectiveness, relative simplicity, and broad applicability across different tasks and modalities. It serves several crucial purposes:
  • Simulating domain shift: By applying transformations that mimic potential variations across different domains (e.g., changes in image intensity, contrast, or noise levels), models can learn to be more robust to these shifts.
  • Addressing data scarcity: In medical imaging, where large, diverse datasets are often challenging to obtain, augmentation can help mitigate the limitations of small sample sizes.
  • Enhancing model generalizability: By exposing the model to a broader range of data variations, it can learn more robust and generalizable features.
For a comprehensive review of general (non-DG) data augmentation methods for MedIA, readers are referred to the survey by Chlap et al[54]. In the following subsections, we focus on augmentation techniques specifically designed or adapted for domain generalization in medical image analysis.
III-A2a   Randomization-based Augmentation
The idea of random augmentation is to generate novel input data by applying random transformations to the original data. Some conventional techniques include randomly applying flipping, rotation, scaling, cropping, adding noise, etc., which are used extensively to improve a model’s generalization performance by reducing overfitting [75]. Li et al[40] developed a novel style transfer network that augments a domain by modifying cardiac images with randomly sampled shape and spatial (i.e., slice index) priors to alleviate the modality-level difference for cardiac segmentation. Liu et al[33] proposed a random amplitude mixup method that randomly mixes the amplitudes of different images for DG for fundus image restoration. Billot et al[76] introduced SynthSeg, a novel approach that leverages domain randomization to train a segmentation network on synthetic brain MRI scans with randomized contrasts and resolutions. This method enables the network to generalize to real scans of varying imaging characteristics without retraining. Similarly, Shen et al[77] proposed RandStainNA, which unifies stain normalization and augmentation techniques by randomly generating virtual stain templates.
III-A2b   Adversarial-based Augmentation
Adversarial-based data augmentation methods operate on the principle of creating adversarial examples that aim to maximize the model’s uncertainty, thereby improving its robustness and generalizability. In this section, we concentrate on data-level adversarial augmentation, while a discussion on model-level adversarial training can be found in Section III-C1c. Tomar et al[43] developed a method that combines knowledge distillation with adversarial-based data augmentation for cross-site medical image segmentation tasks. The process involves the creation of augmented data that is adversarial to the current model, to push the model’s feature representations toward the decision boundary. This is achieved by optimizing and sampling data augmentations that simulate data in the uncertain region of the feature space, thereby improving the model’s ability to generalize from the training data to unseen test data.
III-A2c   Generative Models
@@ III-A2 Data Augmentation
III-A3 Problem-specific Data-level Methods
@@
III-A3a   Cross-modal Generative Models
Cross-modal generative models represent a pioneering paradigm for achieving DG, wherein models are trained to gain knowledge of the data distribution across diverse modalities (e.g., CT, MRI, X-ray, and PET). These models, often based on GANs, generate synthetic data [40] or suitable latent representations [79], which bridge the distributional gap among cross-modalities. This strategy allows us to provide a model especially capable in medical imaging where data could vary greatly due to patient cohorts, hospital practices, or different imaging modalities. As obvious advantages of such a model, it can be highly valuable when one modality is unavailable for a particular patient or when the model is required to generalize to an unseen domain where a different imaging modality is used. Readers are referred to Xie et al[80] for a comprehensive review on cross-modal neuroimage synthesis.
@@
III-B Feature-level Generalization
TABLE V: Summary of feature-level domain generalization methods.